A person wishes to have $28,000 cash for a new car 3 years from now. How much should be placed in an account now, if the account pays 6.1% annual interest rate, compounded weekly?
Please explain to me this question step-by-step.
Homework question from Hans Beauvoir

Answer by mananth(16946)


P = A / (1 + r/n)^nt
P = principal
A = amount accrued
r = rate =6.01% = 0.061
n= number of times compounded. ( weekly) = 52
t= time in years
P = 28,000.00 / (1 + 0.061/52)^((52)(3))
P = 28,000.00 / (1 + 0.001173077)^(156)
Use calculator
P = $23,320.01
The principal investment required to get a total amount of $28,000.00 is $23,320.01